Given that the demand and supply function of commodity X are Qd=20-1/6p while Qs=18+1/6p. Find the equilibrium price and quality for commodity X.
1
Expert's answer
2018-10-12T15:50:09-0400

We can determine the equilibrium price and quantity when Qd = Qs

20-1/6p =18+1/6p

20-18 = 1/6p+1/6p

2=1/3p

p=6

Q=20-1/6 * 6 = 20-1=19